What mass of dilute trioxonitrate(v)acid containing 10% weight per weight of pure acid would be required to desolve 2.5g chalk
1
Expert's answer
2017-01-26T05:09:23-0500
Answer:
20HNO3 + 10CaCO3 = 10Ca(NO3)2+ H2O + 15CO2
M (CaCO3) = 100 g/mol
M (HNO3) = 63 g/mol
n = m/M m = n · M
n (CaCO3) = 2.5/100 = 0.025 mol
n (HNO3) = 2 · n (CaCO3) = 2 · 0.025 = 0.05 mol
m (HNO3) 100% = 0.05 · 63 = 3.15 g
m (HNO3) 10% = 3.15/10 · 100 = 31.5 g